OpenStudy (anonymous):

MEDAL FOR EASY PROBLEM A school band sold 30 raffle tickets, each labeled with a number from 1 to 30. One winning ticket will be drawn. What is the probability that the number of the winning ticket will be a multiple of 4 or the number 19? Express your answer in simplest form.

OpenStudy (anonymous):

i dont get this

OpenStudy (jack1):

30 possible answers 7 tickets = multiples of 4 (4,8,12,16,20,24,28) 1 ticket = 19 therefore: P(multiple of 4 or 19) = 7+1 / 30 = 8/30 = 4/15 ... think that's the simplest fraction you'll get, you can work that out in decimals or % tho if u like?

OpenStudy (jack1):

zat all make sense tho @lordoftanks ?

OpenStudy (anonymous):

(\__/) \. ./ \ /